OpenStudy (anonymous):

Which function below has the largest slope? f(x) x f(x) -3 -10 -1 -7 1 -4 g(x) g(x) over g(x) equals one half x plus 3 h(x) Lance gains 1 level each time he plays his favorite video game. He started at level 10. f(x) g(x) h(x) The functions all have the same slope

OpenStudy (anonymous):

@Hero

OpenStudy (anonymous):

@amistre64 @wio

OpenStudy (anonymous):

The question is unclear. It looks like multiple questions?

OpenStudy (anonymous):

out of the three its wants to know which one has the largest slope? These are the answer choices f(x) g(x) h(x) The functions all have the same slop

ganeshie8 (ganeshie8):

find slope of each function and see which one is largest

ganeshie8 (ganeshie8):

slope of f(x) = ?

ganeshie8 (ganeshie8):

f(x) x f(x) -3 -10 -1 -7 1 -4

ganeshie8 (ganeshie8):

take any two points, and calculate slope using slope formula

OpenStudy (anonymous):

can you help me on that? sorry

ganeshie8 (ganeshie8):

lets take first two points. -3 -10 -1 -7

ganeshie8 (ganeshie8):

-3 -10 x1 y1 -1 -7 x2 y2

ganeshie8 (ganeshie8):

slope = change in y / change in x = \(\large \frac{y_2-y_1}{x_2-x_1}\)

ganeshie8 (ganeshie8):

plugin the values

OpenStudy (anonymous):

oh so -10 -(-7) / -1 - (-3)

ganeshie8 (ganeshie8):

yup !

ganeshie8 (ganeshie8):

simplify

OpenStudy (anonymous):

i got -3 / 2?

ganeshie8 (ganeshie8):

hey wait, u switched y2 and y1, check once

ganeshie8 (ganeshie8):

it shud be -7 -(-10) / -1 - (-3)

ganeshie8 (ganeshie8):

right ?

OpenStudy (anonymous):

so 3/2?

OpenStudy (anonymous):

@ganeshie8

ganeshie8 (ganeshie8):

yup 3/2 = 1.5 slope of f(x) = 1.5 -------------------(1)

ganeshie8 (ganeshie8):

next, find slope of g(x)

ganeshie8 (ganeshie8):

\(\large g(x) = \frac{1}{2} x + 3 \)

ganeshie8 (ganeshie8):

is that how g(x) looks ?

OpenStudy (anonymous):

yes

ganeshie8 (ganeshie8):

u knw how to interpret slope from that equation ?

OpenStudy (anonymous):

not to well

ganeshie8 (ganeshie8):

ok, y = mx+c ^

ganeshie8 (ganeshie8):

m is the slope

OpenStudy (anonymous):

so 1/2 would be my slope?

ganeshie8 (ganeshie8):

Yes :) 1/2 = 0.5 slope of g(x) = 0.5 -------------------(2)

ganeshie8 (ganeshie8):

Lance gains 1 level each time he plays his favorite video game. He started at level 10.

OpenStudy (anonymous):

alright that makes sense

ganeshie8 (ganeshie8):

good :) wat about h(x) ?

ganeshie8 (ganeshie8):

he is gaining 1 level for 1 time he plays so slope = 1/1 = 1

OpenStudy (anonymous):

oh that makes sense....wow im dumb

ganeshie8 (ganeshie8):

its tricky, it ok yuo're very good :)

ganeshie8 (ganeshie8):

so slope of h(x) = 1 ---------------------(3)

ganeshie8 (ganeshie8):

look at (2), (2), (3) once, and make a decision

OpenStudy (anonymous):

so it would be h(x) :)

ganeshie8 (ganeshie8):

careful, look at (1), (2), and (3) lol

OpenStudy (anonymous):

but it wants the largest slope? so it would be h(x)?

ganeshie8 (ganeshie8):

wat about slope of f(x) ?

OpenStudy (anonymous):

OHHHHHHH sorry

ganeshie8 (ganeshie8):

np :)

OpenStudy (anonymous):

thanks for all your help :)

ganeshie8 (ganeshie8):

you wlcme :D